• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name zana

    Log entry time 00:06:12 on June18,2010

    Entry number 327141

    keyword=Swing Shift Summary

    04:00 pm Vacuum pressure at 6 x10^(-5) mbar
    04:35 pm Vacuum pressure at 2.5 x10^(-5) mbar
    04:45 pm Vacuum stable at ~ 2.3 x10^(-5) mbar J.P. Chen called and he is starting the cool down
    07:15 pm Target cooling down completed
    07:18 pm Returning temperature from loop #2 for the target cooling went over 30K, just after had gone down to 24K. Our upper limit is 30K and J.P. Chen suggested to wait a couple of hours to see if the temperature stabilizes
    07:30 pm Called Jack Segal who started to ramping up the magnets for the left and right arm
    10:00 pm magnets for left and right arm seem to be stable
    11:30 pm MCC ready to deliver beam to the Hall. We are trying now to establish 70uA beam with empty target